# MediConfusion: Can you trust your AI radiologist? Probing the reliability of multimodal medical foundation models

### v1

@@ 4.2 Visual prompts in MediConfusion
Free-form visual prompts are intuitive annotations in the input image, such as a red bounding box or an arrow, aimed at highlighting a specific point or area within the image. It is natural to ask whether well-placed visual prompts in medical images, annotated by a doctor, can potentially guide the attention of MLLMs to important areas in the image and thus help provide accurate answers. Such a capability would greatly facilitate human-machine collaboration in healthcare and provide more reliable AI-assisted diagnosis. In the general domain, research has shown that MLLMs typically are unable to efficiently interpret visual prompts without incorporating such task specifically into the training procedure (Cai et al., 2024).
@@ 5.1 Multimodal Large Language Models
Large Language Models (LLMs) such as InstructGPT (Ouyang et al., 2022) and LLaMA (Touvron et al., 2023) have emerged as powerful models capable of performing complex tasks rooted in natural language, including text summarization, coding, and question-answering. LLMs are pretrained on massive text corpora and can be efficiently adapted to downstream tasks. Beyond textual inputs, the LLM pretraining framework has been extended to further modalities, such as images, resulting in multimodal large language models that demonstrate strong visual understanding and reasoning capabilities. LLaVA (Liu et al., 2024) interleaves image representations with the input text of a pretrained LLaMA model and fine-tunes on visual instruction-following data. Flamingo (Alayrac et al., 2022) injects visual information into a frozen LLM via cross-attention. BLIP (Li et al., 2022, 2023) proposes the Q-Former architecture for connecting pretrained vision features to an LLM. More recently, GPT-4o has been trained from scratch on mixed multimodal inputs directly. The resulting MLLMs have achieved great success in visual question answering in the general domain and as multimodal chat agents able to interpret images in the input prompt.
@@
Beyond the general domain, MLLMs are especially promising in automating costly medical tasks, such as analyzing radiology images, generating medical reports or acting as medical conversational agents to provide healthcare advice. There has been substantial research recently to develop medical MLLMs, most often by adapting popular general domain architectures to medical data. Med-Flamingo (Moor et al., 2023) pretrains Flamingo on interleaved image-text medical data sourced from publications and textbooks, unlocking few-shot medical VQA capabilities. Authors of LLaVA-Med (Li et al., 2024) focus on rapid adaptation to the medical domain by fine-tuning LLaVA on filtered image-text pairs from PMC-15M (Zhang et al., 2023a). Authors in Zhang et al. (2023c) generate a large-scale medical VQA dataset from PMC-OA (Lin et al., 2023) which is subsequently used to train MedVInT, a state-of-the-art medical MLLM. Moreover, authors in Wu et al. (2023b) propose a multimodal foundation model for radiology, aligning natural language with 2D and 3D radiology images.
5.2 Encoding visual information in MLLMs
@@
The prevailing approach to incorporate visual information in MLLM training is to leverage contrastive language-image pretrained models as frozen image encoders. CLIP (Radford et al., 2021), and its variants (Cherti et al., 2023), are trained on internet-scale paired image-text data, and thus its representations are readily aligned with natural language, and thus can be effectively combined with language models. The frozen representations are then typically adapted to the feature space of the language model using MLP heads (Liu et al., 2024), Q-Former (Li et al., 2022), cross-attention (Alayrac et al., 2022) or other mechanisms. The image encoder acts as the "eye" of the MLLM as it directly determines what visual information will enter the model. In fact, imperfect compression of relevant visual information is a dominant issue with contemporary MLLMs, resulting in object hallucinations (Li et al., ; Gunjal et al., 2024), fundamental errors in spatial reasoning (Kamath et al., 2023), and inability to understand inter-object relationships (Wu et al., 2024).
@@
As the distribution of general ’internet data’ and medical image-text data is markedly different, CLIP may be unable to capture the intricate structure of medical images with fidelity sufficient for reliable performance. Researchers have proposed CLIP-like models pretrained on large-scale medical data better suited as image encoders for medical MLLMs. LLaVA-Med leverages BiomedCLIP (Zhang et al., 2023b), a foundation model designed for biomedical image-text processing that has been pretrained on PMC-15M. MedVInT uses PMC-CLIP (Lin et al., 2023), a CLIP-style model pretrained on PMC-OA with $1.6M$ medical image-caption pairs. The limitations of image encoders in medical MLLMs have attracted less attention than in the general domain, which is especially troubling due to the safety-critical nature of healthcare applications. Thus, the lack of in-depth understanding of the shortcomings and possible failure modes of the image encoder in the medical MLLM pipeline is an exceedingly pressing concern.
5.3 Medical VQA benchmarks
@@
With the recent rapid advances in developing medical MLLMs, there has been substantial effort in quantifying their performance in a wide range of tasks and areas within the medical domain. VQA-Rad (Lau et al., 2018), SLAKE (Liu et al., 2021), Path-VQA (He et al., 2020) and VQA-Med (Ben Abacha et al., 2021) are widely-used to benchmark the performance of MLLMs in medical VQA. Due to their small size and limited scope, there has been a push for more comprehensive and diverse evaluation datasets. OmniMedVQA (Hu et al., 2024) introduces the largest medical VQA dataset to date, encompassing 12 data modalities and 20 anatomical regions with a total of more than $100k$ images. Authors of Asclepius (Wang et al., 2024b) focus on eliminating data leakage present in other benchmarks and providing human evaluations. GMAI-MMBench (Chen et al., 2024) incorporates problems probing the performance of MLLMs at various perceptual granularities, and targets a well-categorized data structure for ease of preparing customized evaluations. Other benchmarks extend the evaluation task beyond VQA in order to provide a more comprehensive view of model performance. MultiMedEval (Royer et al., 2024) builds a uniform and fair benchmarking framework for multiple tasks including report generation and classification. RadBench (Wu et al., 2023b) focuses on radiology with associated tasks such as modality recognition and disease diagnosis. Authors of CARES (Xia et al., 2024) aim to provide a more holistic view of model performance by focusing on aspects such as fairness, privacy and safety (toxicity) of MLLMs as well as factual correctness (trustfulness).
All of these datasets are aimed at probing the medical knowledge of MLLMs and quantifying their average performance on a wide variety of tasks, modalities and anatomic regions. However, none of these benchmarks are specifically designed to probe the reliability, fundamental limitations and failure modes in the medical domain, all critical aspects in healthcare applications.
@@
Perhaps the closest work in spirit to ours is RadVUQA (Nan et al., 2024), where authors call attention to the critical deficiencies of existing medical MLLMs, revealing a large gap between state-of-the-art MLLMs and clinicians. Their dataset focuses on more fundamental visual question answering and understanding, such as spatial reasoning, anatomic understanding and quantitative reasoning on medical images. However, we go a step further and design a benchmark that stress-tests the visual capabilities of MLLMs by curating questions expected to be challenging for the image processing pipeline of state-of-the-art models.
@@
Related to our work, Tong et al. (2024) has investigated the failure modes of MLLMs originating in ambiguous vision encoding in the general domain. Their study is based on finding CLIP-blind pairs, images that have high similarity in CLIP embedding space, but otherwise have dissimilar low-level image features. However, their methodology is not directly applicable in the medical domain for two reasons. First, CLIP has been pretrained on general domain data and thus it is unable to capture the intricate structure of medical images. Second, their methodology relies on human annotators to describe the difference between a large number of image pairs, which is prohibitively costly in our scenario, as only radiologists are qualified to provide such annotations in the medical setting.
@@ 6 Conclusion
In this paper, we introduce MediConfusion, a challenging medical VQA benchmark designed to probe the limitations of multimodal reasoning in medical MLLMs. In particular, we discover radiology image pairs that, due to ambiguities originating in image encoding, confuse contemporary models despite being dissimilar in the image domain. We leverage an automated pipeline along with the expertise of radiologists to create a dataset of VQA problems that tests the ability of MLLMs to effectively distinguish and answer clinically relevant questions about such confusing pairs. Our benchmark, by construction, cannot be solved by leveraging unimodal priors, and thus, it directly probes multimodal capabilities. We find that existing models achieve performance no better than random guessing on MediConfusion, as models tend to select the same answer option for both images in the pair, raising serious concerns about the reliability of existing MLLMs in a medical setting. In order to guide future research in addressing the limitations of current MLLMs, we identify common failure patterns where models often break and relate them to known limitations in the general domain. We hope that our work sparks further research efforts to improve the reliability of AI for healthcare applications.
@@ A.4 Free Form Evaluation
These scores are the similarities of the MLLM’s answer to the different answer options. If the gap between the higher and lower score is at least $3$, we assign the option with the higher score as the MLLM’s output. Otherwise, we mark the answer as invalid.
@@ B.1 MedFlamingo few-shot prompting
In order for MedFlamingo to produce valid responses, we need to use few-shot prompting. Here, we show three questions and answers from PMC-VQA benchmarks Zhang et al. (2023c). The following is the prompt we used for MC evaluation:

### v2

@@ 4.2 Visual prompts in MediConfusion
Free-form visual prompts are intuitive annotations in the input image, such as a red bounding box or an arrow, aimed at highlighting a specific point or area within the image. It is natural to ask whether well-placed visual prompts in medical images, annotated by a doctor, can potentially guide the attention of MLLMs to important areas in the image and thus help provide accurate answers. Such a capability would greatly facilitate human-machine collaboration in healthcare and provide more reliable AI-assisted diagnosis. In the general domain, research has shown that MLLMs typically are unable to efficiently interpret visual prompts without incorporating such task specifically into the training procedure [3].
@@ 5 Related Work
Multimodal Large Language Models – Large Language Models (LLMs) such as InstructGPT [29] and LLaMA [37] have emerged as powerful models capable of performing complex tasks rooted in natural language, including text summarization, coding, and question-answering. LLMs are pretrained on massive text corpora and can be efficiently adapted to downstream tasks. Beyond textual inputs, the LLM pretraining framework has been extended to further modalities, such as images, resulting in multimodal large language models that demonstrate strong visual understanding and reasoning capabilities. LLaVA [25] interleaves image representations with the input text of a pretrained LLaMA model and fine-tunes on visual instruction-following data. Flamingo [1] injects visual information into a frozen LLM via cross-attention. BLIP [19, 20] proposes the Q-Former architecture for connecting pretrained vision features to an LLM. More recently, GPT-4o has been trained from scratch on mixed multimodal inputs directly.
@@
Beyond the general domain, MLLMs are especially promising in automating costly medical tasks, such as analyzing radiology images, generating medical reports or acting as medical conversational agents to provide healthcare advice. There has been substantial research recently to develop medical MLLMs, most often by adapting popular general domain architectures to medical data. Med-Flamingo [27] pretrains Flamingo on interleaved image-text medical data sourced from publications and textbooks, unlocking few-shot medical VQA capabilities. Authors of LLaVA-Med [18] focus on rapid adaptation to the medical domain by fine-tuning LLaVA on filtered image-text pairs from PMC-15M [49]. Authors in Zhang et al. [51] generate a large-scale medical VQA dataset from PMC-OA [22] which is subsequently used to train a medical MLLM. Moreover, authors in Wu et al. [42] propose a multimodal foundation model, RadFM, for radiology, aligning natural language with 2D and 3D radiology images.
@@
Encoding visual information in MLLMs – The prevailing approach to incorporate visual information in MLLM training leverages contrastive language-image pretrained models as frozen image encoders. CLIP [31], and its variants [5], are trained on internet-scale paired image-text data, and thus its representations are readily aligned with natural language, and can be effectively combined with language models. The frozen representations are then adapted to the feature space of the language model using MLP heads [25], Q-Former [19], cross-attention [1] or other mechanisms. The image encoder acts as the "eye" of the MLLM as it directly determines what visual information enter the model. In fact, imperfect compression of relevant visual information is a dominant issue with contemporary MLLMs, resulting in object hallucinations [21, 11], fundamental errors in spatial reasoning [15], and inability to understand inter-object relationships [43].
@@
As the distribution of general ’internet data’ and medical image-text data is markedly different, CLIP may be unable to capture the intricate structure of medical images with fidelity sufficient for reliable performance. Researchers have proposed CLIP-like models pretrained on large-scale medical data better suited as image encoders for medical MLLMs. LLaVA-Med leverages BiomedCLIP [50], a foundation model designed for biomedical image-text processing that has been pretrained on PMC-15M. MedVInT uses PMC-CLIP [22], a CLIP-style model pretrained on PMC-OA with $1.6M$ medical image-caption pairs. The limitations of image encoders in medical MLLMs have attracted less attention than in the general domain, which is especially troubling due to the safety-critical nature of healthcare applications. Thus, the lack of in-depth understanding of the shortcomings and possible failure modes of the image encoder in the medical MLLM pipeline is an exceedingly pressing concern.
@@
Medical VQA benchmarks – With the recent rapid advances in developing medical MLLMs, there has been substantial effort in quantifying their performance in a wide range of tasks and areas within the medical domain. VQA-Rad [16], SLAKE [23], Path-VQA [12] and VQA-Med [2] are widely-used to benchmark the performance of MLLMs in medical VQA. Due to their small size and limited scope, there has been a push for more comprehensive and diverse evaluation datasets. OmniMedVQA [13] introduces the largest medical VQA dataset to date, encompassing 12 data modalities and 20 anatomical regions with a total of more than $100k$ images. Authors of Asclepius [39] focus on eliminating data leakage present in other benchmarks and providing human evaluations. GMAI-MMBench [4] incorporates problems probing the performance of MLLMs at various perceptual granularities, and targets a well-categorized data structure for ease of preparing customized evaluations. Other benchmarks extend the evaluation task beyond VQA in order to provide a more comprehensive view of model performance. MultiMedEval [32] builds a uniform and fair benchmarking framework for multiple tasks including report generation and classification. Micro-Bench [26] evaluates the visual understanding capabilities of MLLMs across diverse microscopy modalities. RadBench [42] focuses on radiology with associated tasks such as modality recognition and disease diagnosis. Authors of CARES [45] aim to provide a more holistic view of model performance by focusing on aspects such as fairness, privacy and safety of MLLMs as well as factual correctness.
@@
All of these datasets are aimed at probing the medical knowledge of MLLMs and quantifying their average performance on a wide variety of tasks, modalities and anatomic regions. However, none of these benchmarks are specifically designed to probe the reliability, fundamental limitations and failure modes in the medical domain, all critical aspects in healthcare applications. Perhaps the closest work in spirit to ours is RadVUQA [28], where authors call attention to the critical deficiencies of existing medical MLLMs, revealing a large gap between state-of-the-art MLLMs and clinicians. Their dataset focuses on more fundamental visual question answering and understanding, such as spatial reasoning, anatomic understanding and quantitative reasoning on medical images. However, we go a step further and design a benchmark that stress-tests the visual capabilities of MLLMs by curating questions expected to be challenging for their image processing pipeline.
@@
Related to our work, Tong et al. [36] has investigated the failure modes of MLLMs originating in ambiguous vision encoding in the general domain. Their study is based on finding CLIP-blind pairs, images that have high similarity in CLIP embedding space, but otherwise have dissimilar low-level image features. However, their methodology is not directly applicable in the medical domain for two reasons. First, CLIP has been pretrained on general domain data and thus it is unable to capture the intricate structure of medical images. Second, their methodology relies on human annotators to describe the difference between a large number of image pairs, which is prohibitively costly in our scenario, as only radiologists are qualified to provide such annotations in the medical setting.
@@ 6 Conclusion
In this paper, we introduce MediConfusion, a challenging medical VQA benchmark designed to probe the limitations of multimodal reasoning in medical MLLMs. In particular, we discover radiology image pairs that, due to ambiguities originating in their multimodal embedding spaces, confuse contemporary models despite being dissimilar in the image domain. We leverage an automated pipeline along with the expertise of radiologists to create a dataset of VQA problems that tests the ability of MLLMs to effectively distinguish and answer clinically relevant questions about such confusing pairs. Our benchmark, by construction, cannot be solved by leveraging unimodal priors, and thus, it directly probes multimodal capabilities. We find that most existing models achieve performance no better than random guessing on MediConfusion, as models tend to select the same answer option for both images in the pair, raising serious concerns about the reliability of existing MLLMs in a medical setting. In order to guide future research in addressing the limitations of current MLLMs, we identify common failure patterns where models often break and relate them to known limitations in the general domain. We hope that our work sparks further research efforts to improve the reliability of AI for healthcare applications.
@@ A.4 Free Form Evaluation
These scores are the similarities of the MLLM’s answer to the different answer options. If the gap between the higher and lower score is at least $k=3$, which we call the score gap threshold, we assign the option with the higher score as the MLLM’s output. Otherwise, we mark the answer as invalid.
@@ B.1 MedFlamingo few-shot prompting
In order for MedFlamingo to produce valid responses, we need to use few-shot prompting. Here, we show three questions and answers from PMC-VQA benchmarks [51]. The following is the prompt we used for MC evaluation:
@@
[…]
Model Prompt 0 Prompt 1 Prompt 2 Prompt 3 Prompt 4 Prompt 5 Prompt 6 Prompt 7 Prompt 8 Prompt 9 Prompt 10 All Prompts
GPT-4o $19.32\pm 1.53$ $19.43\pm 0.66$ $18.01\pm 0.95$ $17.33\pm 0.96$ $18.30\pm 0.80$ $19.66\pm 0.81$ $19.43\pm 1.62$ $20.40\pm 1.12$ $19.38\pm 1.15$ $19.43\pm 1.50$ $19.66\pm 1.11$ $19.12\pm 0.04$
InstructBLIP $11.65\pm 1.42$ $5.57\pm 0.91$ $3.75\pm 0.77$ $4.38\pm 1.48$ $2.67\pm 1.19$ $1.31\pm 0.57$ $4.26\pm 0.99$ $4.43\pm 1.86$ $1.42\pm 0.77$ $12.05\pm 3.15$ $1.25\pm 0.66$ $4.79\pm 0.12$
LLaVA-Med $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Appendix C Ablation on prompting
In this section, we perform an ablation study on the effect of prompt variations on model performance. We test GPT-4o, InstructBLIP, and LLaVA-Med with MC evaluation. We manually create 10 prompt variations different from the original, shown in Table 5. We do not rely on LLMs for rephrasing the original prompt to avoid biases and artifacts originating in LLM-generated text. To capture variability in accuracy both due to prompting and stochasticity, we sample $10$ answers for each model, and for each prompt. Table 4 shows our results.
First, we observe that the performance of GPT-4o is robust to the prompt format, but the performance is still below random guessing. InstructBLIP is more sensitive to the particular input prompt, as demonstrated by significantly reduced performance on some prompts. Lastly, LLaVA-Med struggles with the MC evaluation, as it has not been specifically trained to follow the multiple-choice format, and prompt engineering does not fix this issue.
Table 5: List of prompts we use to evaluate the sensitivity of model performance to prompt variations. The second column shows the high-level rationale for the variation, and the specific prompt is included in the last column.
# Change Prompt
0 Original Based on the image, choose the correct option for the following question.
Question:
A:
B:
Answer with the option’s letter from the given choices directly. Your answer should be just one letter.
Answer:
1 Shorten Answer the following question about the image.
Question:
A:
B:
Answer with the option’s letter directly.
Answer:
2 Move answer options to the end Answer the following question about the image with the correct option’s letter directly.
Question:
A:
B:
Answer:
3 Instruct the model for extra care and attention Carefully and skillfully review the image and answer the following question.
Question:
A:
B:
Answer with the option’s letter directly.
Answer:
4 Define AI expert role You are an expert radiologist AI with deep knowledge in radiology image analysis. Based on the image, choose the correct option for the following question.
Question:
A:
B:
Answer with the option’s letter directly.
Answer:
5 Rephrase the instruction to answer with a single letter Based on the image, choose the correct option for the following question.
Question:
A:
B:
Simply respond with ’A’ or ’B’ based on your answer.
Answer:
6 Ask the model nicely Based on the image, please choose the correct option for the following question.
Question:
A:
B:
Please answer with the option’s letter from the given choices directly.
Answer:
7 Remove new lines from formatting Based on the image, choose the correct option for the following question. Question: Option A: Option B: Answer with the option’s letter from the given choices directly. Your answer should be just one letter. Answer:
8 Give more explanation You are given a radiology image, based on which you will need to carefully answer a medical question related to the image. You will be given two answer options: A and B from which you have to choose the right answer. The correct answer can always be determined just by looking at the radiology image itself. Take a close look at the image, think carefully, and answer the following question.
Question:
A:
B:
Answer with the option’s letter from the given choices directly. Your answer should be just one letter.
Answer:
9 No instructions Question:
A:
B:
Correct answer option:
10 Emphasize importance You are given a radiology image, and your critical task is to take a close look and choose the correct option for the following question. It is extremely important to answer the following question correctly.
Question:
A:
B:
Answer with the option’s letter from the given choices directly. Your answer should be just one letter.
Answer:
Appendix D Fine-tuning on MediConfusion
In this section, we investigate the effect of fine-tuning medical MLLMs on MediConfusion. Fine-tuning on the specific benchmark typically helps the model to better understand the format of the VQA and can greatly boost the performance of the model on the particular benchmark.
We create train and test splits from MediConfusion by randomly sampling $50\%$ of confusing pairs from each category to train on, and hold out the other half for evaluation. We ensure that the two images from a confusing pair always belong to the same split in order to avoid leakage. As each pair may belong to multiple categories, the above process results in two splits of slightly different size ($84$ pairs in train and $92$ pairs in test), but with category distribution close to the full dataset. We following common procedure to fine-tune LLaVA-Med on the train split, involving fine-tuning the language model and the multimodal adapter, keeping the image encoder frozen. We perform a search over learning rates within $\left[10^{-6},5\cdot 10^{-6},10^{-5},10^{-4},2\cdot 10^{-4}\right]$ and select $5\cdot 10^{-6}$. We keep other fine-tuning hyperparameters at their recommended values. We perform full fine-tuning, as opposed to parameter-efficient fine-tuning approaches, and train for up to $1000$ epochs. Train and test (set) accuracy across epochs are depicted in Figure 6.
Interestingly, we observe that the model struggles to achieve $100\%$ set accuracy on the training set, even after a very high number of training epochs (typical fine-tuning is within $1-50$ epochs), further supporting our hypothesis that the model is unable to differentiate between image pairs due to fundamental ambiguity in their embeddings. In particular, the vision features appear so similar to the model that even fine-tuning the whole language model is insufficient. This experiment suggests that MediConfusion can only be solved by improving the vision encoder directly.
Furthermore, we find that fine-tuning helps LLaVA-Med learn the multiple-choice format (set accuracy on the test split increased from $0\%$ to around $20\%$). However, its performance is still well below random guessing, as the knowledge learned on the training set does not generalize to the test set.
Figure 6: Evolution of train and test set accuracies during LLaVA-Med fine-tuning over epochs.
Appendix E Ablation study on dataset difficulty
[…]
MediConfusion-Easy MediConfusion-Unfiltered MediConfusion
Model Set acc. Confusion Indiv. acc. Set acc. Confusion Indiv. acc. Set acc. Confusion Indiv. acc.
GPT-4o 98.38 1.32 99.04 18.01 77.83 55.97 18.75 75.00 56.25
LLaVA-Med 34.21 65.59 66.90 4.86 95.75 51.06 0.00 97.16 49.72
We investigate the effect of the difficulty of confusing pairs in our benchmark construction. Specifically, we create an ’easy’ version of MediConfusion from image pairs that look dissimilar to medical vision-language models, which we refer to as MediConfusion-easy. We extract pairs from ROCO that have low BiomedCLIP similarity ($sim_{med}\leq 0.4$). As radiologist feedback is prohibitively costly and time-consuming, we build MediConfusion-easy by following the same automatic pipeline as we use in curating MediConfusion, but skip radiologist feedback. Thus, MediConfusion-easy may include questions that are clinically irrelevant or have incorrect answers. The resulting dataset consists of $988$ pairs ($1976$ VQA problems).
To make sure that the performance gap between MediConfusion and MediConfusion-easy is not due to the lack of radiologist feedback, we also create MediConfusion-unfiltered using identical hyperparameters and procedure to the original benchmark curation, but without radiologist feedback. We match the number of samples to MediConfusion-easy.
We evaluate GPT-4o (MC evaluation) and LLaVA-Med (PS evaluation) on MediConfusion-easy. Table 6 summarizes our results. First, we highlight that removing radiologist feedback does not necessarily improve performance (GPT-4o has slightly lower set accuracy, LLaVA-Med performed somewhat better on unfiltered data). Furthermore, our experiment demonstrates that models can indeed differentiate between image pairs in our confusing pair-based VQA format, if the images are different enough, as evidenced by the nearly $100\%$ accuracy of GPT-4o on MediConfusion-easy. Thus, this experiment further supports the need for a strict selection criteria for confusing pairs, as easier variants of our VQA may be unable to identify the shortcomings of state-of-the-art models.
We believe that selecting challenging pairs with high BiomedCLIP similarity, as done in our work, contributes to the difficulty in $2$ ways. First, due to the ambiguity of input vision features, the model is unable to resolve enough details in the medical images that is sufficient to answer the questions correctly. Second, the questions themselves are also more challenging, as the difference between hard confusing image pairs is more nuanced. Thus, the model needs more medical knowledge to correctly answer the question, even if the vision embeddings are sufficiently dissimilar. We hypothesize that this medical knowledge is lacking in the case of the Gemini model in our experiments, where we observe poor performance despite low confusion. To highlight how questions are more general and potentially require less medical knowledge, we show an example from MediConfusion-easy in 7, where identifying the anatomical region is sufficient for solving the VQA pair.
Figure 7: A sample VQA pair from MediConfusion-easy. Identifying the anatomical region is sufficient to answer the pair correctly.
Appendix F Investigating feature similarity of vision encoders
We find confusing pairs based on BiomedCLIP similarities to construct MediConfusion. The resulting poor performance across all models implies that they struggle differentiating between the confusing pairs, however it is unclear how similar the images are in the respective models’ own image feature space. In this section, we analyze the similarity of MediConfusion pairs in the feature space of vision encoders used by open-source models in our experiments: CLIP and the visual encoder of RadFM. RadFM has its own vision encoder trained on medical images, including 3D radiology images. Figure 8 shows the histogram of similarities between confusing pairs based on the cosine similarity of BiomedCLIP, CLIP, and RadFM visual embeddings. We average pool across the token dimension in the output of vision encoders to produce the visual embedding.
As these plots show, the images have high similarity in not only BiomedCLIP’s, but other vision encoders’ embedding spaces, both trained on general-domain data and medical data.
Figure 8: Histograms of the cosine similarity of confusing pairs in different visual representations.
Appendix G Reversing the role of DINO and BiomedCLIP
As an ablation study, we explore samples that have high DINO similarity but low BiomedCLIP similarity. In particular, we searched for image pairs for which $sim_{med}\leq 0.5$ and $sim_{gen}\geq 0.5$. Figure 9 shows some of these samples.
Figure 9: Sample with high DINO similarity but low BiomedCLIP similarity
Appendix H Exploring Techniques to Improve Performance on MediConfusion
In this section, we examine various methods we employed to improve model performance on MediConfusion and analyze their effectiveness.
H.1 Mixture of Features
Following [36], we investigate the effect of incorporating DINOv2 embeddings into the image encoding of LLaVA-Med. Specifically, we concatenate the DINOv2 embeddings with the existing CLIP embeddings to create an enriched image representation. We then apply the feature alignment and instruction tuning procedures outlined in [24] and train the model from scratch. Additionally, we follow the medical-specific feature alignment and instruction tuning methodologies from [18].
Table 7 compares the performance of this modified model with the original LLaVA-Med. While we observe a slight improvement in accuracy, overall performance remains poor, indicating that richer visual features alone are insufficient to solve MediConfusion.
Table 7: Performance of LLaVA-Med model with added DINOv2 embeddings
[…]
Method MC GD PS MC GD PS MC GD PS Set acc. Indiv. acc.
LLaVA-Med 0.00 0.00 1.14 23.58 49.72 49.72 100.00 99.43 97.16 1.14 49.72
LLaVA-Med + DINOv2 embed. 2.84 2.84 2.27 45.45 50.57 50.00 95.54 95.45 95.45 2.84 50.57
H.2 Best-of-$N$
The best-of-$N$ method is a widely used approach for enhancing MLLM responses([33], [10], [47]). In this technique, we generate $N$ different responses and select the best one based on a predefined scoring metric. For this study, we use GPT-4o and explore different scoring strategies.
H.2.1 Oracle-based Scoring
To establish an upper bound on model performance, we employ an oracle-based scoring method, where the oracle has access to the ground truth solutions. We hypothesize that if the model is capable of correctly answering the questions with some reasonable, albeit low, probability, then improved prompting strategies can potentially elicit its medical knowledge. Specifically, we follow the FF evaluation framework, where we first present only the question to the model and generate a response. We then provide GPT-4o with the question, answer choices, and the generated response, asking it to determine which option is closest to the generated answer. We sample $N=100$ responses for each question. Furthermore, we lower the score gap threshold to $k=2$. We then measure the frequency with which the correct response appears among these generations. We accept an answer if the frequency of the correct answer exceeds a certain threshold. Table 8 presents the results of this experiment. As we can see, even under these lenient conditions, the model performs poorly, with its performance declining significantly as the acceptance threshold increases.
Table 8: GPT-4o Best of 100 performance with an oracle as the scoring metric
[…]
1 47.16 68.18
2 37.50 62.50
3 32.95 59.38
4 27.84 55.97
5 24.43 53.41
H.2.2 Self-Ranking
Now, we consider a more realistic setup in which we do not have access to the ground truth. Similar to the previous method, we generate $N$ different samples, then we show the question, options, and all of the generated answers to GPT-4o and ask it to select the best response. We then pass the selected answer as the final answer of the model. Table 9 provides the results of this section for different choices of $N$. Note that since we are using the FF evaluation, in which the model cannot see the answer options prior to generating the response, the performance is much lower compared to the MC evaluation results in Table 1. These results show that even when we are aggregating $50$ responses, the model’s performance remains poor, suggesting the presence of fundamental limitations that prevent it from effectively solving these problems.
Table 9: GPT-4o best-of-$N$ performance with a judge as the scoring metric
[…]
10 5.11 33.24 81.58 41.48
20 4.55 30.97 81.33 43.75
50 7.39 36.08 79.76 38.64
H.2.3 Majority voting
Self-consistency [40] sampling can improve reasoning performance by marginalizing over multiple reasoning paths to aggregate the final answer. We generate $N$ different answers using the MC evaluation and then choose the most frequent one as the model’s final answer. As shown in Table 10, with $N=10$ we observe $5\%$ improvement, but increasing $N$ further yields negligible gains.
Table 10: GPT-4o Majority voting with $N$ answers
[…]
1 18.75 56.25 75.00
10 23.30 57.39 68.00
100 22.73 57.39 69.32